OpenStudy (anonymous):

Use the quadratic formula to solve the equation. If necessary, round to the nearest hundredth x² + 3 = –4x

sam (.sam.):

x^(2)+3=-4x x^(2)+4x+3=0 --------------------------------------- a=1 b=4 c=3 \[\huge x=\frac{-b \pm \sqrt{b ^{2}-4ac}}{2a}\] \[\huge x=\frac{-(4) \pm \sqrt{16-4(3)}}{2}\] --------------------------------------- \[\huge x=\frac{-(4)+2}{2}\] OR \[\huge x=\frac{-(4)-2}{2}\] x=-3,-1